# Deep reinforcement learning-based active flow control of an elliptical cylinder: transitioning from an elliptical cylinder to a circular cylinder and a flat plate

### v2

@@ III.2 Effectiveness of AFC
The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers. Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction. In our research, the control strategy derived from the DRL-based AFC system dynamically responds in real time to variations in the flow state. This emphasizes its capability to adaptively adjust to changing conditions, demonstrating the robustness and responsiveness of the AFC framework. Furthermore, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
@@ III.3 Analysis of the Controlled Flow
Figure 13: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed. The streamwise snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively.
@@
At $Ar=1$ and $0.75$, during the $T_{1}$ phase, a standard Bénard–von Kármán vortex street forms behind the elliptical cylinder, characterized by a series of regularly spaced vortices. These vortices alternately shed from both sides of the cylinder, creating a regular pattern of vortex arrangement. Moving into the second phase ($T_{2}$), similar to cases with higher aspect ratios ($Ar=1.75$, 1.5, 1.25), the recirculation bubbles behind the cylinder are elongated longitudinally, and the originally shed vortices are also elongated and narrowed, with shearing interactions occurring between them. By the third time point ($T_{3}$), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability.
@@
Figure 14: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed. The streamwise snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
@@
Fig. 14 delineates scenarios where the suppression of vortex shedding from elliptical cylinders is not entirely effective for $Ar$ of 0.5, 0.25, 0.1, and 0. At $T_{1}$, the vortex shedding behind the cylinder is fully developed, $T_{2}$ represents the control phase, and $T_{3}$ provides a snapshot of the flow field after control cessation. For $Ar=0.5$ and $Ar=0.25$, During the $T_{1}$ phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake. In the $T_{2}$ phase, as control commences, the recirculation bubble is noted to elongate longitudinally to some extent, with the recirculation area appearing relatively symmetric and stable. The intervention of synthetic jets delays the detachment of vortices, which are stretched and deformed due to shear with the side walls. By $T_{3}$, the initially generated vortices are significantly elongated and deformed, exiting the computational domain, yet the recirculation bubble behind the cylinder remains unstable and continues to oscillate, leading to new vortex shedding. The emergence of new irregular vortex shedding highlights persistent instability behind the elliptical cylinder. For elliptical cylinders with $Ar=0.5$ and $Ar=0.25$, the DRL-based control strategy does not completely suppress vortex shedding. The instability of the wake field is mitigated, demonstrating partial efficacy of the control strategy in manipulating flow dynamics, although full stability is not achieved.
@@
For $Ar=0.1$ and $Ar=0$, during $T_{1}$, the vortex shedding frequency and intensity from the cylinder are more severe compared to $Ar=0.5$, with enhanced flow instability. Multi-level vortex shedding occurs, with primary and secondary vortices differing in shape, frequency, and intensity. At T2, vortices behind the cylinder are elongated and delayed, and the recirculation area significantly enlarges. The instability of the recirculation area manifests as asymmetry, swinging, and oscillation phenomena. By $T_{3}$, although the original shedding vortices are elongated and expelled from the computational domain, new, unstable, and irregular shedding occurs. Shed vortices swinging to the side walls are sheared, stretched, and undergo deformation, resulting in irregularly shaped vortices. This highlights the persistent challenges in managing flow dynamics and instability at lower $Ar$.
@@ IV CONCLUSIONS
In this study, we investigate the use of DRL-based AFC to suppress vortex shedding, reduce drag, and mitigate lift fluctuations of elliptical cylinders at a Reynolds number of 100. The $Ar$ of the elliptical cylinder ranges from an ellipse ($Ar=2.0$) to a circle ($Ar=1.0$), and finally to a flat plate ($Ar=0$). We employ the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder. The primary objective of this work is to gain further insights into the control capabilities of DRL algorithms for complex flow systems. The main findings of this study can be summarized as follows:
@@
  • At $Ar=2$, the baseline flow is already in a stable state. When AFC is applied, the agent does not perform any actions that would disturb the flow field. This indicates that the DRL agent is capable of recognizing a stable flow condition and making intelligent decisions accordingly. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%. These data indicate that the AFC strategy based on DRL is capable of significantly minimizing the periodically fluctuating drag coefficient to minimal values and stabilizing it, while also maintaining the oscillating lift coefficient near zero. This demonstrates the efficacy of DRL in AFC applications, effectively reducing the lift and drag around elliptical cylinder with $Ar$ ranging from 0.75 to 1.75. Visualizations of velocity fields reveal that the original periodically shedding vortices transition to stable flow fields with no vortex shedding after flow control. Instantaneous vorticity maps during the control process show that with the activation of synthetic jets, the vortices downstream of the elliptical cylinder are progressively elongated. The originally unstable recirculation area behind the cylinder transforms into a stable, symmetrical recirculation bubble, ceasing the production of new vortices. These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), and its capability of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder. Moreover, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively. This underscores that the flow control strategy efficiently utilizes less than 1% of external excitation energy, exemplifying its energy-saving capability for effective flow control.
@@
  • When the $Ar$ of the elliptical cylinder decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the drag coefficient and lift coefficient are 26.9%, 35.9%, 37.6%, and 43.6%, respectively, while the corresponding reduction rates for the lift coefficient are 68.0%, 66.1%, 50.2%, and 60.1%. Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, vortex shedding phenomena have not yet stabilized compared to cases with larger $Ar$ values. The initial phase of control exhibits distinct fluctuations, followed by low-frequency and low-amplitude oscillations in the drag and lift coefficients, which are lower than the baseline. Particularly, for $Ar$ values of 0.1 and 0, the drag and lift coefficients experience a sharp decrease initially, followed by moderate-amplitude, low-frequency periodic oscillations. This indicates that for elliptical cylinders with $Ar$ ranging from 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag. From the vorticity plots of the control process, for $Ar$ values of 0.5 and 0.25, the recirculation region behind the elliptical cylinder is elongated during the initial phase of control, with symmetric recirculation bubbles. As the control progresses, the recirculation region becomes unstable, exhibiting lateral oscillations and the shedding of long, narrow vortices. Similarly, for $Ar$ values of 0.1 and 0, the recirculation region undergoes elongation during the control phase, with high-frequency and large-amplitude oscillations, accompanied by the shedding of new vortices. After a significant period of control, the shedding of vortices is somewhat alleviated, with a decrease in both the frequency and intensity of vortex shedding. Correspondingly, for elliptical cylinders with $Ar$ values of 0.5, 0.25, 0.1, and 0, the ratios of the mass flow rates of the synthetic jets relative to the incoming flow rate are 4.9%, 12.6%, 49.9%, and 52.1%, respectively. This suggests that despite the use of higher external energy, the lift coefficient cannot be completely controlled to prevent oscillations and fluctuations.
@@
In conclusion, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy. The control policies based on the PPO algorithm significantly reduce lift around the elliptical cylinder, greatly diminish drag, and completely suppress vortex shedding in the wake. Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in lift and drag coefficients. These control results demonstrate the adaptability of DRL-based AFC strategies to various geometric shapes, their robustness against uncertainties, and their ability to handle complex problems. This makes DRL-based AFC particularly suitable for controlling complex fluid dynamic systems.

### v3

@@ III.2 Effectiveness of AFC
Furthermore, in Tab. 2, we normalize the mass flow rates of the synthetic jets against the incoming far-field flow rate, reflecting their relative strength compared to the incoming flow. $\overline{a}$ ratio (%) represents the ratio of the mass flow rate of the synthetic jets to the flow rate from the far-field inflow boundary. For $Ar$ greater than 0.75, the ratio of synthetic jets to the inflow is less than 1%, indicating that the mass flow rate utilized in the flow control strategy proposed by the PPO agent is minimal, in line with the energy efficiency requirements of AFC technologies. For $Ar$ ranging from 0.75 to 1.75, the elliptical cylinder achieves a reduction in $C_{D}$ and attenuates up to 95% of $C_{L}$ with very low energy consumption from external excitation. However, as $Ar$ decreases from 0.5 to 0.1, the ratio of synthetic jets to the inflow increases significantly. Despite this increase in energy input, the capability to attenuate $C_{L}$ decreases. This indicates that at lower $Ar$ values, the instability in the wake field cannot be fully controlled, even with higher energy from external excitation. The control strategies derived from DRL-based AFC technology exhibit greater complexity compared to conventional harmonic forcing implemented in previous works (Bergmann, Cordier, and BrancherBergmann, Cordier, and Brancher (2005)), highlighting the value of employing ANN as controllers. Typically, approximately 93% of the drag induced by the shedding of vortices is suppressed by the discovered control laws, with minimal jet intensities required for drag reduction. In our research, the control strategy derived from the DRL-based AFC system dynamically responds in real time to variations in the flow state. This emphasizes its capability to adaptively adjust to changing conditions, demonstrating the robustness and responsiveness of the AFC framework. Furthermore, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the magnitude of the mass flow rate injected by the jets normalized by the mass flow rate of the mainstream intersecting with the cylinder is within 1%.
The above content details the effectiveness of the flow control strategy, with the corresponding DRL training process curves for elliptical cylinders with different $Ar$ values presented in Fig. 13. First, we observed the overall trend of the reward function. For $Ar$ values ranging from 1.75 to 0, we noted that the reward function increases with the number of training iterations, eventually converging to a stable value. However, for $Ar=2$, after 5,000 training iterations, the reward function reaches a plateau but remains below zero, indicating that although the reward function has converged, the agent is unable to achieve the desired objective. Next, we examined the convergence performance of the reward function. For $Ar$ values ranging from 1.75 to 0, the reward function exhibits clear convergence points, stabilizing after a plateau phase. This indicates that the agent has learned effective strategies, achieving stable returns under these conditions. However, the convergence speed of DRL training varies significantly across different $Ar$ values. For $Ar$ values of 1.75, 1.5, and 1.25, convergence signs appear around 2,000 iterations. For $Ar$ values of 1 and 0.75, convergence signs emerge around 1,000 iterations. Despite setting the total training iterations to 5,000 for $Ar$ values between 1.75 and 0.75, we observed that the reward function converges much earlier for $Ar$ values of 1 and 0.75, indicating a faster convergence speed and the agent’s ability to quickly learn effective strategies. For $Ar=0.5$, convergence signs appear after 4,000 iterations. For $Ar=0.25$, convergence signs emerge after 8,000 iterations. For $Ar=0.1$, convergence signs are observed after 10,000 iterations. For $Ar=0$, convergence signs appear after 12,000 iterations. In the range of $Ar=0.5$ to 0, the DRL training convergence speed is significantly slower compared to $Ar$ values greater than 0.5. This reflects a lower training efficiency and greater difficulty in exploring effective strategies for the agent when $Ar$ is less than or equal to 0.5. Furthermore, in the case of different $Ar$, DRL training was conducted in parallel using 60 CFD environments, allowing the DRL algorithm to collect data simultaneously from multiple independent simulations. The parameters of the ANN were updated every 60 iterations, which means that the agent and the environment only interacted without updating the policy network for the first 60 episodes. After 60 episodes, the ANN network parameters started to update, resulting in an initial learning curve that exhibits a noticeable step-like increase. This parallel acceleration effect has also been observed in the research conducted by Rabault and KuhnleRabault and Kuhnle (2019a) and Wang and XuWang and Xu (2024a).
@@ III.3 Analysis of the Controlled Flow
Figure 14: When $Ar=2$, 1.75, 1.5, 1.25, 1, 0.75, the vortex shedding of the elliptical cylinder can be completely suppressed. The instantaneous vorticity snapshots of the wake flow field behind an elliptical cylinder, where the vortex shedding process is suppressed, feature key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These represent the instances of the onset of control (when the vortex is fully developed), during the control process, and at the result of control, respectively.
@@
At $Ar=1$ and $0.75$, during the $T_{1}$ phase, a standard Bénard–von Kármán vortex street forms behind the elliptical cylinder, characterized by a series of regularly spaced vortices. These vortices alternately shed from both sides of the cylinder, creating a regular pattern of vortex arrangement. Moving into the second phase ($T_{2}$), similar to cases with higher aspect ratios ($Ar=1.75$, 1.5, 1.25), the recirculation bubbles behind the cylinder are elongated longitudinally, and the originally shed vortices are also elongated and narrowed, with shearing interactions occurring between them. By the third time point ($T_{3}$), the vortices behind the elliptical cylinder are completely suppressed, and the regions of positive and negative vorticity in the flow field nearly achieve symmetry, demonstrating high stability. Fig. 15 delineates scenarios where the suppression of vortex shedding from elliptical cylinders is not entirely effective for $Ar$ of 0.5, 0.25, 0.1, and 0. At $T_{1}$, the vortex shedding behind the cylinder is fully developed, $T_{2}$ represents the control phase, and $T_{3}$ provides a snapshot of the flow field after control cessation.
@@
Figure 15: When $Ar=0.5$, 0.25, 0.1 and 0, vortex shedding from the elliptical cylinder is not entirely suppressed. The instantaneous vorticity snapshots of the wake flow field behind the elliptical cylinder capture key moments designated as $T_{1}$, $T_{2}$, and $T_{3}$. These correspond to the initiation of control (when the vortex is fully developed), the early phase of the control process, the mid-to-late phase of the control process, and the termination phase of control, respectively.
@@
For $Ar=0.5$ and $Ar=0.25$, During the $T_{1}$ phase, vortices are observed detaching from the cylinder with certain frequency and intensity, forming a Kármán vortex street in the downstream wake. In the $T_{2}$ phase, as control commences, the recirculation bubble is noted to elongate longitudinally to some extent, with the recirculation area appearing relatively symmetric and stable. The intervention of synthetic jets delays the detachment of vortices, which are stretched and deformed due to shear with the side walls. By $T_{3}$, the initially generated vortices are significantly elongated and deformed, exiting the computational domain, yet the recirculation bubble behind the cylinder remains unstable and continues to oscillate, leading to new vortex shedding. The emergence of new irregular vortex shedding highlights persistent instability behind the elliptical cylinder. For elliptical cylinders with $Ar=0.5$ and $Ar=0.25$, the DRL-based control strategy does not completely suppress vortex shedding. The instability of the wake field is mitigated, demonstrating partial efficacy of the control strategy in manipulating flow dynamics, although full stability is not achieved.
@@
For $Ar=0.1$ and $Ar=0$, during $T_{1}$, the vortex shedding frequency and intensity from the cylinder are more severe compared to $Ar=0.5$, with enhanced flow instability. Multi-level vortex shedding occurs, with primary and secondary vortices differing in shape, frequency, and intensity. At $T_{2}$, vortices behind the cylinder are elongated and delayed, and the recirculation area significantly enlarges. The instability of the recirculation area manifests as asymmetry, swinging, and oscillation phenomena. By $T_{3}$, although the original shedding vortices are elongated and expelled from the computational domain, new, unstable, and irregular shedding occurs. Shed vortices swinging to the side walls are sheared, stretched, and undergo deformation, resulting in irregularly shaped vortices. This highlights the persistent challenges in managing flow dynamics and instability at lower $Ar$.
@@ IV CONCLUSIONS
In this study, we investigated the use of DRL-based AFC to suppress vortex shedding, reduce drag, and mitigate lift fluctuations of elliptical cylinders at a $Re=100$. The $Ar$ of the elliptical cylinder ranged from an elliptical cylinder ($Ar=2.0$) to a circle cylinder ($Ar=1.0$), and finally to a flat plate ($Ar=0$). We employed the PPO algorithm to precisely control the mass flow rates of synthetic jets located on the upper and lower surfaces of the elliptical cylinder. The primary objective of this work was to gain further insights into the control capabilities of DRL algorithms for complex flow systems. The main findings of this study can be summarized as follows:
@@
  • At $Ar=2$, the baseline flow is already in a stable state. When AFC is applied, the agent does not perform any actions that would disturb the flow field. This indicates that the DRL agent is capable of recognizing a stable flow condition and making intelligent decisions accordingly. However, for $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the reduction rates in the drag coefficient around the cylinder are respectively 0.9%, 2.1%, 3.9%, 8.0%, and 15.7%, while the reduction rates in lift coefficient are 99.7%, 99.4%, 98.9%, 98.9%, and 95.2%. The originally periodic oscillations of the drag coefficient are reduced to their minimum value and stabilized. Simultaneously, over 95% of the lift coefficient is suppressed, ultimately stabilizing close to zero. With the activation of synthetic jets, the oscillating recirculation bubble in the baseline flow is gradually elongated and transformed into a stable and symmetrical recirculation bubble. Simultaneously, the existing vortices in the wake region become progressively elongated, ceasing the generation of new vortices. The originally periodic shedding of vortices transitions into a stable flow field, eliminating the phenomenon of vortex shedding. These results underscore the robust adaptability of DRL-based AFC across different elliptical cylinder shapes ($Ar$ between 0.75 and 1.75), and its capability of learning optimal control strategies directly through interaction with the environment to achieve a reduction in drag and lift coefficients around the cylinder. Moreover, for elliptical cylinders with $Ar$ of 1.75, 1.5, 1.25, 1, and 0.75, the mass flow rate ratios of synthetic jets relative to the inlet flow rate are 0.7%, 0.9%, 0.6%, 0.1%, and 1.0% respectively. This indicates the effectiveness of the flow control strategy, which utilizes less than 1% of external excitation energy, showcasing its significant energy-saving attribute.
@@
  • When the $Ar$ of the elliptical cylinder decreases to 0.5, 0.25, 0.1, and 0, the reduction rates of the drag coefficient and lift coefficient are 26.9%, 35.9%, 37.6%, and 43.6%, respectively, while the corresponding reduction rates for the lift coefficient are 68.0%, 66.1%, 50.2%, and 60.1%. Despite the effective reduction in both drag and lift coefficients achieved by the DRL-based AFC strategy, vortex shedding phenomena have not yet stabilized compared to cases with larger $Ar$ values. The drag and lift coefficients exhibit significant fluctuations during the initial stages of control, eventually settling into oscillatory states with smaller amplitudes and frequencies compared to the baseline flow. In particular, for elliptical cylinders with $Ar$ values of 0.1 and 0, the lift and drag coefficients continue to exhibit periodic oscillations. This indicates that for elliptical cylinders within the $Ar$ range of 0.5 to 0, the AFC control strategy can partially suppress lift and reduce drag. For $Ar$ of 0.5 and 0.25, the recirculation region behind the elliptical cylinder initially elongates, exhibiting symmetric recirculation bubbles. As control progresses, the recirculation region becomes unstable, displaying lateral oscillations and the shedding of long, narrow vortices. For $Ar$ values of 0.1 and 0, the recirculation region remains unstable throughout the control phase, accompanied by the shedding of new vortices. However, compared to the baseline flow, the frequency and intensity of vortex shedding decrease. Correspondingly, for elliptical cylinders with $Ar$ values of 0.5, 0.25, 0.1, and 0, the ratios of the mass flow rates of the synthetic jets relative to the incoming flow rate are 4.9%, 12.6%, 49.9%, and 52.1%, respectively. This indicates that despite using higher external energy, it is still not possible to suppress the oscillations and fluctuations in the lift coefficient.
@@
To summarize, for elliptical cylinders with $Ar$ ranging from 1.75 to 0.75, intelligent agents trained using the PPO algorithm have adeptly learned to adopt an energy-efficient flow control strategy. The control policies based on the PPO algorithm significantly reduce $C_{L}$ around the elliptical cylinder, greatly diminish $C_{D}$, and completely suppress vortex shedding in the wake. Regarding elliptical cylinder with $Ar$ values ranging from 0.5 to 0, the control strategies obtained from agents trained with the PPO algorithm can moderately alleviate vortex shedding and reduce the amplitude and frequency of fluctuations in $C_{D}$ and $C_{L}$. These control results demonstrate the adaptability of DRL-based AFC strategies to various geometric shapes, their robustness against uncertainties, and their ability to handle complex problems.